OpenStudy (clara1223):

Solve the inequality for x given that (x-6)/(x-3)(x-9)<0 a) (-inf,-6)U(6,9) b) (-inf,3)U(6,inf) c) (-inf,6)U(6,inf) d) (-inf,-3)U(6,9) e) (-inf,3)U(6,9)

OpenStudy (anonymous):

this is not as hard as it looks changes sign at 3,6,9

OpenStudy (anonymous):

divide number line up in to four intervals pick some interval and see if it is positive or negative there there, then you will know all

OpenStudy (clara1223):

3,6,9 would all be open circles right?

OpenStudy (anonymous):

yeah but you don't need no circles, you need intervals

OpenStudy (anonymous):

\[(-\infty, 3),(3,6),(6,9),(9\infty)\] are your intervals check on one then alternate

OpenStudy (clara1223):

somehow i got that (-inf,3)U(3,6)U(9,inf)

OpenStudy (clara1223):

I have no idea what im doing wrong

OpenStudy (clara1223):

oh wait is it e?

OpenStudy (anonymous):

you have two adjacent intervals \((-\infty,3)\) and \((3,6)\) that can't be right

OpenStudy (anonymous):

it changes sign at 3 so if it is positive on one interval it must be negative on the other did you check on any one interval? with some practice you won't have too, but it doesn't hurt

OpenStudy (clara1223):

i did a little math wrong and forgot that i'm looking for negatives, not positives

OpenStudy (anonymous):

what interval did you check?

OpenStudy (clara1223):

i got that (-inf,3) was negative, (3,6) was positive, (6,9) was negative, and (9,inf) was positive

OpenStudy (anonymous):

i would also point out that one only of the above answers is possible, even if you didn't check anything

OpenStudy (anonymous):

yeah at the risk of repeating myself you only have to check one if you know it is negative on \((-\infty,3)\) then it must be \[\huge -,+,-,+\]

OpenStudy (clara1223):

so the correct answer is (-inf,3)(6,9), right?

OpenStudy (anonymous):

and yeah it is E

jimthompson5910 (jim_thompson5910):

The graph confirms the answer you got https://www.desmos.com/calculator/ipckz8kxhy the portion on the interval (-inf, 3) is below the x axis the portion on the interval (6,9) is below the x axis everything else is either on the x axis or above it